# Recursive identification with regularization and on-line hyperparameters estimation

### v1

@@ Abstract
This paper presents a regularized recursive identification algorithm with simultaneous on-line estimation of both the model parameters and the algorithms hyperparameters. A new kernel is proposed to facilitate the algorithm development. The performance of this novel scheme is compared with that of the recursive least-squares algorithm in simulation.
@@ 1 Introduction
It is only since since the beginning of the 2010s, that the Bayesian approach has attracted a considerable interest in the identification field with the emergence of kernel-based identification schemes, coming from Machine Learning (Pillonetto and de Nicolau (2010)). By applying some regularization techniques, namely modifying a least squares criterion, the bias-variance trade-off can be significantly improved, especially if the initial problem is ill conditioned (Pillonetto et al. (2014)). The most significant kernels embedding a prior knowledge of intrinsic systems properties (such as stability and smoothness of the impulse response) are the Stable-Spline (SS) Kernel (Pillonetto and de Nicolau (2010)), the Diagonal Correlated (DC) and the Tuned-Correlated (TC) kernels (Chen et al. (2012)). These kernels are tuned with a number of hyperparameters whose values can be determined beforehand from data, using a Hierarchical Bayes approach (see Pillonetto et al. (2022) p. 107). So far, to the authors knowledge, this kernel approach has been employed in a recursive context in (Prando et al. (2016)), (Romeres et al. (2016)) and (Illg et al. (2022)).
In the sequel one presents a novel recursive and regularized identification scheme for the estimation of the impulse response of a discrete Linear Time Invariant (LTI) system, and where the model is a Finite Impulse Response (FIR) filter. The proposed scheme is more easily implementable online than the one disclosed in (Romeres et al. (2016)). Like in this last reference it is assumed that the prior is not known at initialization, which requires to estimate online the prior hyperparameters at the same time as the parameters. In general, in regularized schemes, the number of Kernel hyperparameters is reduced (see Pillonetto et al. (2022)). Here, in order to make the algorithm implementable on line, one proposes unusually a prior with a number of hyperparameters equal to the number of parameters, the prior structure (smoothness of the impulse response and stability) being imposed by putting some constraints to those hyperparameters. This prior is finally very close to the TC prior (Chen et al. (2012)). The online estimation of the hyperparameters aims at minimizing the cost of the marginal likelihood criterion (Empirical Bayes approach) (Pillonetto et al. (2022), chap.4), by using a recursive gradient with projection, while the parameters are estimated with an algorithm quite similar to RLS.
The paper is structured as follows: Section 2 shows how the recursive least-squares can be implemented in case of regularized criterion, following a Bayesian approach, Section 3 proposes a prior structure inspired from the TC prior, Section 4 details the algorithm mixing on-line parameters and hyperparameters estimation. At last Section 5 displays some simulations showing the performances of this novel algorithm.
@@ 2.1 Classical propagation equations for recursive estimation
Our purpose is to estimate the vector $\theta$ where
@@
$J_{0}(t)=\sum_{i=1}^{t}\left(y(i)-\phi^{T}(i-1)\hat{\theta}(i)\right)^{2}+\hat{\theta}^{T}(t)\sigma^{2}\Pi_{0}^{-1}\hat{\theta}(t),$ (3)
$\hat{\theta}(t)$ being the estimation of $\theta$ at time $t$. The positive definite matrix $\Pi_{0}$ is the regularization matrix (corresponding to the kernel functions in a Machine Learning approach). By cancelling out the criterion gradient with respect to the vector $\hat{\theta}$, one finds immediately
@@
In a Bayesian perspective, the prior knowldege is incorporated in $\Pi_{0}$ (which corresponds to the covariance matrix of this prior). Therefore, in a recursive perspective, incorporating this prior is equivalent to specify $F(0)$. However, when one performs an identification in real-time, in general one cannot use some available data so as to specify this prior (by applying for example the Hierarchical Bayes approach) allowing to determine the optimal prior hyperparameters, by maximizing a marginal likelihood function. It is the reason why, we consider in this paper the issue of computing on-line the regularisation matrix $\Pi(t)$ in parallel with the computation of $\hat{\theta}(t)$.
@@ 2.2 Case where the regularization matrix is no longer invariant
Let us assume that by a specific estimation algorithm (which will be detailed below), the regularization matrix $\Pi$ is updated at each sample time, yielding $\Pi(t)$. We assume that $\Pi$ is expressed from a vector of hyperparameters $\eta$ which is no longer constant. From (5) and (2.1), one can write now
@@
The propagation equations of $F^{-1}$ becomes
@@
On the other hand, from (2.1), one has
@@ 2.3 Marginal Likelihood in a recursive context
In the hierarchical Bayes approach, the vector of hyperparameters $\eta$ is estimated from data at first, by maximizing the Marginal Likelihood function $L(\eta|Y)$. Afterwards the maximum a posteriori (MAP) estimate $\hat{\theta}=argmax\;\textbf{p}(\theta|Y)$ is computed, p being the probability density function (pdf).
In a recursive context, where data are obtained on-line, and available up to time $t+1$, the maximization of the Marginal Likelihood function $L(\eta(t+1)|Y(t+1))$ and the computation of the MAP $\hat{\theta}(t+1)=argmax\;\textbf{p}(\theta(t+1)|Y(t+1),\eta(t+1))$ must be computed alternatively, so as to take into account only available data at time $t+1$. For this purpose set
@@
The derivative of $\log L(\eta(t+1)|Y(t+1))$ with respect to the $k$ entry of $\eta$ (denoted as $\eta_{k}$) is
$\frac{\partial log(L(t+1))}{\partial\eta_{k}}=\frac{1}{2}Tr\left(\left[\sigma^{-2}\left(\Pi(t+1)-\right.\right.\right.\\ \left.\left.\left.-\hat{\theta}(t+1)\hat{\theta}^{T}(t+1)\right)-F(t+1)\right]\frac{\partial\sigma^{2}\Pi^{-1}(t+1)}{\partial\eta_{k}}\right)$ (15)
@@ 3.1 Prior structure
where
$U=\begin{bmatrix}1&1&\cdots&1\\ 0&1&\cdots&1\\ 0&0&\ddots&\vdots\\ 0&\cdots&0&1\end{bmatrix}$ (17)
and
${W(t,\eta)=\begin{bmatrix}e^{\eta_{1}(t)}&0&\cdots&0\\ 0&e^{\eta_{2}(t)}&\cdots&0\\ \vdots&\cdots&\ddots&\vdots\\ 0&0&\cdots&e^{\eta_{n}(t)}\end{bmatrix}}$ (18)
with $\eta^{T}(t)=[\eta_{1}(t)\;\eta_{2}(t)\;\cdots\eta_{n}(t)]$.
@@
$W_{i,i}(t,\eta)=\lambda(t)W_{i-1,i-1}(t,\eta)$
with
$10<\lambda(t)<$ 1
This prior structure differs from the TC kernel only because of the term $W_{n,n}$ (see Carli et al. (2017)). One can now perform a change of basis where the parameters estimate is $\hat{\theta}^{{}^{\prime}}$ and the regressor $\phi^{{}^{\prime}}(t)$ such that
${\hat{\theta}^{{}^{\prime}}(t)=U^{-1}\hat{\theta}(t)}$ (19)
${\phi^{{}^{\prime}}(t)=U^{T}\phi(t)}$ (20)
${F^{{}^{\prime}}(t+1,\eta(t+1))=U^{-1}F(t+1,\eta(t+1))U^{-T}=}\\ {\sum_{i=0}^{t}\phi^{{}^{\prime}}(i)\phi^{{}^{\prime}T}(i)+\sigma^{2}W(t+1,\eta(t+1))}$ (21)
@@
${\hat{\theta}^{{}^{\prime}}(t+1,\eta(t+1))=}\\ {\left(\sum_{i=1}^{t}\phi^{{}^{\prime}}(i)\phi^{{}^{\prime}}(i)^{T}+\sigma^{2}W^{-1}(\eta(t+1))\right)^{-1}\sum_{i=1}^{t}\phi^{{}^{\prime}}(i)y(i)}$ (22)
Thus, in the basis where $\hat{\theta}^{{}^{\prime}}$ is the estimated parameters vector and $\phi^{{}^{\prime}}$ the regressor, the associated prior $W$ is nothing but a kernel matrix $K^{{}^{\prime}}$ corresponding to the DI Kernel (see Pillonetto et al. (2022), p.145), with
@@
The prior $\Pi(t,\eta)$ as defined in (16), (17), and (18) under the constraints C1 and C2 is stable.
@@
The prior $\Pi(t,\eta)$ as defined in (16), (17), and (18) has the property of maximum entropy.
Proof: See Appendix B.
@@ 3.2 On-line hyperparameters estimation
The estimation procedure of the vector $\eta$ consists in finding $\eta$ that minimizes $J_{h}(t)=\log(L(\eta(t)|Y(t)))$. By combining (15), (19)–(21), one can write (the dependence with respect to $t$ being omitted for sake of simplicity)
@@ 4.1 Update of the adaptation gain an estimated parameters vector
$F^{{}^{\prime}-1}(t+1,\eta(t+1))-F^{{}^{\prime}-1}(t+1,\eta(t))=\\ =-\sigma^{2}\gamma\left(\operatorname{diag}\left(\Psi_{p}(t)\right)\right)^{-1}.$ (31)
@@
${F^{{}^{\prime}-1}(t+1,\eta(t+1))-F^{{}^{\prime}-1}(t+1,\eta(t))=}\\ {\sigma^{2}\gamma F^{{}^{\prime}}(t+1,\eta(t))\left(diag\left(\Psi_{p}(t)\right)\right)^{-1}F^{{}^{\prime}}(t+1,\eta(t))}$ (32)
@@
$\hat{\theta}^{{}^{\prime}}(t+1,\eta(t+1))$ $=F^{{}^{\prime}}(t+1,\eta(t+1))\sum_{i=0}^{t}\phi^{{}^{\prime}}(t)y(i+1)$ (33a)
$\hat{\theta}^{{}^{\prime}}(t+1,\eta(t))$ $=F^{{}^{\prime}}(t+1,\eta(t))\sum_{i=0}^{t}\phi^{{}^{\prime}}(t)y(i+1)$ (33b)
and by combining (31) and (4.1), one gets
${\hat{\theta}^{{}^{\prime}}(t+1,\eta(t+1))=}\\ {\left(I_{n}+\sigma^{2}\gamma F^{{}^{\prime}}(t+1,\eta(t+1))\left(diag\left(\Psi_{p}(t)\right)\right)^{-1}\right)\hat{\theta}(t+1,\eta(t))}$ (34)
@@ 4.2 Summary
The algorithm computing alternatively the estimation of parameters and hyperparameters is given by
${\hat{\theta}^{\prime}(t+1,\eta(t))=}\\ {\hat{\theta}^{\prime}(t,\eta(t))+F^{\prime}(t,\eta(t))\phi^{\prime}(t)\frac{\varepsilon^{o}(t+1)}{1+\phi^{{}^{\prime}T}(t)F^{\prime}(t,\eta(t))\phi^{{}^{\prime}}(t)}}$ (35)
${F^{{}^{\prime}}(t+1,\eta(t))=}\\ {F^{{}^{\prime}}(t,\eta(t))-\frac{F^{{}^{\prime}}(t,\eta(t))\phi^{\prime}(t)\phi^{{}^{\prime}T}(t)F^{\prime}(t,\eta(t))}{1+\phi^{{}^{\prime}T}F^{\prime}(t,\eta(t))\phi^{\prime}(t)}}$ (36)
${\eta(t+1)=\eta(t)-\gamma\Psi_{p}(t)}$ (37)
${F^{{}^{\prime}-1}(t+1,\eta(t+1))=F^{{}^{\prime}-1}(t+1,\eta(t))+}\\ {\sigma^{2}\gamma F^{{}^{\prime}}(t+1,\eta(t))\left(diag\left(\Psi_{p}(t)\right)\right)^{-1}F^{{}^{\prime}}(t+1,\eta(t))}$ (38)
${\hat{\theta}^{{}^{\prime}}(t+1,\eta(t+1))=}\\ {\left(I_{n}+\sigma^{2}\gamma F^{{}^{\prime}}(t+1,\eta(t+1))\left(\operatorname{diag}\left(\Psi_{p}(t)\right)\right)^{-1}\right)\hat{\theta}(t+1,\eta(t))}$ (39)
@@ Appendix A Proof of Lemma 1
$\frac{\partial log\left|\Sigma^{-1}(t+1)\right|}{\partial\eta_{k}}=\\ =Tr\left[\left(\sigma^{-2}\Pi(t+1)-F(t+1)\right)\frac{\sigma^{2}\partial\Pi^{-1}(t+1)}{\partial\eta_{k}}\right]$ (46)
@@
$\frac{\partial l(t+1)}{\partial\eta_{k}}=-\sigma^{-2}Y^{T}(t+1)\cdot\\ \cdot\Phi\frac{\partial\left(\sigma^{2}\Pi^{-1}(t+1)+\Phi^{T}(t)\Phi(t)\right)^{-1}}{\partial\eta_{k}}\Phi^{T}(t)Y(t+1)$ (48)
@@
$\frac{\partial l(t+1)}{\partial\eta_{k}}=\sigma^{-2}\sum_{i=1}^{t}y(i+1)\phi^{T}(i)F(t+1)\frac{\partial\sigma^{2}\Pi^{-1}}{\partial\eta_{k}}F(t+1)\cdot\\ \cdot\sum_{i=1}^{t}\phi(i)y(i+1)$ (49)
@@
$\frac{\partial l(t+1)}{\partial\eta_{k}}=\sigma^{-2}\hat{\theta}^{T}(t+1)\frac{\partial\sigma^{2}\Pi^{-1}}{\partial\eta_{k}}\hat{\theta}(t+1)$ (50)
@@
$\frac{\partial log(L(t+1))}{\partial\eta_{k}}=\frac{1}{2}\left(Tr\left[\left(\sigma^{-2}\Pi(t+1)-F(t+1)\right)\right.\right.\\ \left.\left.\frac{\partial\sigma^{2}\Pi^{-1}(t+1)}{\partial\eta_{k}}\right]-\frac{1}{\sigma^{2}}\hat{\theta}^{T}(t+1)\frac{\partial\sigma^{2}\Pi^{-1}(t+1)}{\partial\eta_{k}}\hat{\theta}(t+1)\right)$ (51)

### v2

@@ Abstract
This paper presents a regularized recursive identification algorithm with simultaneous on-line estimation of both the model parameters and the algorithms hyperparameters. A new kernel is proposed to facilitate the algorithm development. The performance of this novel scheme is compared with that of the recursive least squares algorithm in simulation.
@@ 1 Introduction
It is only since the beginning of the 2010s, that the Bayesian approach has attracted a considerable interest in the identification field with the emergence of kernel-based identification schemes, coming from Machine Learning (Pillonetto and de Nicolau (2010)). By applying some regularization techniques, namely modifying a least squares criterion, the bias-variance trade-off can be significantly improved, especially if the initial problem is ill conditioned (Pillonetto et al. (2014)). The most significant kernels embedding a prior knowledge of intrinsic systems properties (such as stability and smoothness of the impulse response) are the Stable-Spline (SS) Kernel (Pillonetto and de Nicolau (2010)), the Diagonal Correlated (DC) and the Tuned-Correlated (TC) kernels (Chen et al. (2012)). These kernels are tuned with a number of hyperparameters whose values can be determined beforehand from data, using a Hierarchical Bayes approach (see (Pillonetto et al., 2022, p. 107)). So far, to the authors knowledge, this kernel approach has been employed in a recursive context in (Prando et al. (2016)), (Romeres et al. (2016)) and (Illg et al. (2022)).
In the sequel one presents a novel recursive and regularized identification scheme for the estimation of the impulse response of a discrete Linear Time Invariant (LTI) system, and where the model is a Finite Impulse Response (FIR) filter. The proposed scheme is more easily implementable online than the one disclosed in (Romeres et al. (2016)). Like in this last reference, it is assumed that the prior is not known at initialization, which requires to estimate online the prior hyperparameters at the same time as the parameters. In general, in regularized schemes, the number of Kernel hyperparameters is reduced (see Pillonetto et al. (2022)). Here, in order to make the algorithm implementable on line, one proposes unusually a prior with a number of hyperparameters equal to the number of parameters, the prior structure (smoothness of the impulse response and stability) being imposed by putting some constraints to those hyperparameters. This prior is finally very close to the TC prior (Chen et al. (2012)). The online estimation of the hyperparameters aims at minimizing the cost of the marginal likelihood criterion (Empirical Bayes approach) (Pillonetto et al., 2022, chap.4), by using a recursive gradient with projection, while the parameters are estimated with an algorithm quite similar to recursive least squares (RLS).
The paper is structured as follows: Section 2 shows how the recursive least-squares can be implemented in case of regularized criterion, following a Bayesian approach, Section 3 proposes a prior structure inspired from the TC prior, Section 4 details the algorithm mixing on-line parameters and hyperparameters estimation. At last, Section 5 displays some simulations showing the performances of this novel algorithm.
@@ 2.1 Classical propagation equations for recursive estimation
Our purpose is to estimate the vector $\theta$, where
@@
$J_{0}(t)=\sum_{i=1}^{t}\left(y(i)-\phi^{T}(i-1)\hat{\theta}(t)\right)^{2}+\hat{\theta}^{T}(t)\sigma^{2}\Pi_{0}^{-1}\hat{\theta}(t),$
$\hat{\theta}(t)$ being the estimation of $\theta$ at time $t$, and
$\phi(t)=[u(t-1),~{}u(t-2),\ldots u(t-n)]^{T}$
The positive definite matrix $\Pi_{0}$ is the regularization matrix (corresponding to the kernel functions in a Machine Learning approach). By cancelling out the criterion gradient with respect to the vector $\hat{\theta}$, one finds immediately
@@
In a Bayesian perspective, the prior knowldege is incorporated in $\Pi_{0}$ (which corresponds to the covariance matrix of this prior). Therefore, in a recursive estimation algorithm, incorporating this prior is equivalent to specify $F(0)$. However, when one performs an identification in real-time, in general one cannot use some available data so as to specify this prior (by applying for example the Hierarchical Bayes approach) allowing to determine the optimal prior hyperparameters, by maximizing a marginal likelihood function. It is the reason why, we consider in this paper the issue of computing on-line the regularisation matrix $\Pi(t)$ in parallel with the computation of $\hat{\theta}(t)$.
@@ 2.2 Case where the regularization matrix is no longer invariant
Let us assume that by a specific estimation algorithm (which will be detailed below), the regularization matrix $\Pi$ is updated at each sample time, yielding $\Pi(t)$. We assume that $\Pi$ is expressed from a vector of hyperparameters $\eta$ which is no longer constant. From (4) and (5), one can write now
@@
The propagation equation of $F^{-1}$ becomes
@@
On the other hand, from (7), one has
@@ 2.3 Marginal Likelihood in a recursive context
In the hierarchical Bayes approach, the vector of hyperparameters $\eta$ is estimated from data at first, by maximizing the Marginal Likelihood function $L(\eta|Y)$. Afterwards the maximum a posteriori (MAP) estimate $\hat{\theta}=argmax\;\textbf{p}(\theta|Y)$ is computed, p being the probability density function (pdf). In a recursive context, where data are obtained on-line, and available up to time $t+1$, the maximization of the Marginal Likelihood function $L(\eta(t+1)|Y(t+1))$ and the computation of the MAP $\hat{\theta}(t+1)=argmax\;\textbf{p}(\theta(t+1)|Y(t+1),\eta(t+1))$ must be computed alternatively, so as to take into account only available data at time $t+1$. For this purpose set
@@
The derivative of $\log L(\eta(t+1)|Y(t+1))$ with respect to the $k$ entry of $\eta(t+1)$ (denoted as $\eta_{k}(t+1)$) is
$\frac{\partial log(L(t+1))}{\partial\eta_{k}(t+1)}=\frac{1}{2}Tr\left(\left[\left(\Pi(t+1)-\right.\right.\right.\\ \left.\left.\left.-\hat{\theta}(t+1)\hat{\theta}^{T}(t+1)\right)-\sigma^{2}F(t+1)\right]\frac{\partial\Pi^{-1}(t+1)}{\partial\eta_{k}(t+1)}\right)$ (14)
@@ 3.1 Prior structure
where $U$ is upper triangular with every nonzero element equal to one and $W(t,\eta)=\operatorname{diag}(\exp(\eta(t)))$, with $\eta^{T}(t)=[\eta_{1}(t)\;\eta_{2}(t)\;\cdots\eta_{n}(t)]$.
@@
$W_{i,i}(t,\eta)=\exp(-\alpha(t))W_{i-1,i-1}(t,\eta)$
with $0<\alpha(t)$. This prior structure differs from the TC kernel only because of the term $W_{n,n}$ (see (Carli et al., 2017, eqs. (20)–(22))). As the TC kernel, the proposed one is well suited to damped systems. One can now perform a change of basis where the parameters estimate is $\hat{\theta}^{{}^{\prime}}$ and the regressor $\phi^{{}^{\prime}}(t)$ such that
${\hat{\theta}^{{}^{\prime}}(t)=U^{-1}\hat{\theta}(t)},\quad{\phi^{{}^{\prime}}(t)=U^{T}\phi(t)},$ (16)
${F^{{}^{\prime}-1}(t+1,\eta(t+1))=U^{-1}F^{-1}(t+1,\eta(t+1))U^{-T}=}\\ {\sum_{i=0}^{t}\phi^{{}^{\prime}}(i)\phi^{{}^{\prime}T}(i)+\sigma^{2}W^{-1}(t+1,\eta(t+1))}$ (17)
@@
${\hat{\theta}^{{}^{\prime}}(t+1,\eta(t+1))=}\\ {\left(\sum_{i=0}^{t}\phi^{{}^{\prime}}(i)\phi^{{}^{\prime}}(i)^{T}+\sigma^{2}W^{-1}(\eta(t+1))\right)^{-1}\sum_{i=0}^{t}\phi^{{}^{\prime}}(i)y(i)}$ (18)
Thus, in the basis where $\hat{\theta}^{{}^{\prime}}$ is the estimated parameters vector and $\phi^{{}^{\prime}}$ the regressor, the associated prior $W$ is nothing but a kernel matrix $K^{{}^{\prime}}$ corresponding to the DI Kernel (see Chen et al. (2012)), with
@@
The prior $\Pi(t,\eta)$ as defined in (15) under the constraints C1 and C2 is stable.
@@
The prior $\Pi(t,\eta)$ as defined in (15) has the property of maximum entropy.
The proof follows (Carli, 2014, Section IV) and is omitted.
@@ 3.2 On-line hyperparameters estimation
The estimation procedure of the vector $\eta$ consists in finding $\eta$ that minimizes $J_{h}(t)=-\log(L(\eta(t)|Y(t)))$. By combining (14), (16)–(17), one can write (the dependence with respect to $t$ being omitted for sake of simplicity)
@@ 4.1 Update of the adaptation gain an estimated parameters vector
$F^{{}^{\prime}-1}(t+1,\eta(t+1))-F^{{}^{\prime}-1}(t+1,\eta(t))=\\ =\sigma^{2}\left(W^{-1}(\eta(t+1))-W^{-1}(\eta(t))\right).$ (27)
@@
$F^{{}^{\prime}}(t+1,\eta(t+1))=F^{{}^{\prime}}(t+1,\eta(t))-\sigma^{2}F^{{}^{\prime}}(t+1,\eta(t))\cdot\\ \cdot\left(W^{-1}(\eta(t+1))-W^{-1}(\eta(t))\right)F^{{}^{\prime}}(t+1,\eta(t))$ (28)
@@
$\hat{\theta}^{{}^{\prime}}(t+1,\eta(t+1))$ $=F^{{}^{\prime}}(t+1,\eta(t+1))\sum_{i=0}^{t}\phi^{{}^{\prime}}(t)y(i+1)$
$\hat{\theta}^{{}^{\prime}}(t+1,\eta(t))$ $=F^{{}^{\prime}}(t+1,\eta(t))\sum_{i=0}^{t}\phi^{{}^{\prime}}(t)y(i+1)$
and by combining with (28), one gets
$\hat{\theta}^{{}^{\prime}}(t+1,\eta(t+1))=\left(I_{n}-\sigma^{2}F^{{}^{\prime}}(t+1,\eta(t))\cdot\right.\\ \left.\cdot\left(W^{-1}(\eta(t+1))-W^{-1}(\eta(t))\right)\right)\hat{\theta}(t+1,\eta(t))$ (29)
@@ 4.2 Summary
The algorithm computing alternatively the estimation of parameters and hyperparameters is given by [^1] Remark: since there is no explicit matrix inversion in the update of $F$, the algorithm’s computational complexity is $O(n^{2})$ at each sampling instant. Note also that the matrix inversion in (13) has complexity $O(t^{3})$.
$\hat{\theta}^{\prime}(t+1,\eta(t))=\hat{\theta}^{\prime}(t,\eta(t))+\\ +F^{\prime}(t,\eta(t))\phi^{\prime}(t)\frac{\varepsilon^{o}(t+1)}{1+\phi^{{}^{\prime}T}(t)F^{\prime}(t,\eta(t))\phi^{{}^{\prime}}(t)}$ (30a)
$F^{{}^{\prime}}(t+1,\eta(t))=F^{{}^{\prime}}(t,\eta(t))-\\ -\frac{F^{{}^{\prime}}(t,\eta(t))\phi^{\prime}(t)\phi^{{}^{\prime}T}(t)F^{\prime}(t,\eta(t))}{1+\phi^{{}^{\prime}T}F^{\prime}(t,\eta(t))\phi^{\prime}(t)}$ (30b)
${\eta(t+1)=\eta(t)-\gamma\Psi_{p}(t)}$ (30c)
$F^{{}^{\prime}}(t+1,\eta(t+1))=F^{{}^{\prime}}(t+1,\eta(t))-\sigma^{2}F^{{}^{\prime}}(t+1,\eta(t))\cdot\\ \cdot\left(W^{-1}(\eta(t+1))-W^{-1}(\eta(t))\right)F^{{}^{\prime}}(t+1,\eta(t))$ (30d)
$\hat{\theta}^{{}^{\prime}}(t+1,\eta(t+1))=\left(I_{n}-\sigma^{2}F^{{}^{\prime}}(t+1,\eta(t))\cdot\right.\\ \left.\cdot\left(W^{-1}(\eta(t+1)-W^{-1}(\eta(t))\right)\right)\hat{\theta}(t+1,\eta(t))$ (30e)
@@ Appendix A Proof of Lemma 1
$\frac{\partial log\left|\Sigma^{-1}(t+1)\right|}{\partial\eta_{k}(t+1)}=\\ =Tr\left[\left(\sigma^{-2}\Pi(t+1)-F(t+1)\right)\frac{\sigma^{2}\partial\Pi^{-1}(t+1)}{\partial\eta_{k}(t+1)}\right]$ (37)
@@
$\frac{\partial l(t+1)}{\partial\eta_{k}(t+1)}=-\sigma^{-2}Y^{T}(t+1)\cdot\\ \cdot\Phi\frac{\partial\left(\sigma^{2}\Pi^{-1}(t+1)+\Phi^{T}(t)\Phi(t)\right)^{-1}}{\partial\eta_{k}(t+1)}\Phi^{T}(t)Y(t+1)$ (39)
@@
$\frac{\partial l(t+1)}{\partial\eta_{k}(t+1)}=\sigma^{-2}\sum_{i=1}^{t}y(i+1)\phi^{T}(i)F(t+1)\cdot\\ \cdot\frac{\partial\sigma^{2}\Pi^{-1}}{\partial\eta_{k}(t+1)}F(t+1)\sum_{i=1}^{t}\phi(i)y(i+1)$ (40)
@@
$\frac{\partial l(t+1)}{\partial\eta_{k}(t+1)}=\sigma^{-2}\hat{\theta}^{T}(t+1)\frac{\partial\sigma^{2}\Pi^{-1}}{\partial\eta_{k}(t+1)}\hat{\theta}(t+1)$ (41)
@@
$\frac{\partial log(L(t+1))}{\partial\eta_{k}(t+1)}=\frac{1}{2}\left(Tr\left[\left(\sigma^{-2}\Pi(t+1)-F(t+1)\right)\right.\right.\\ \left.\left.\frac{\partial\sigma^{2}\Pi^{-1}(t+1)}{\partial\eta_{k}(t+1)}\right]-\frac{1}{\sigma^{2}}\hat{\theta}^{T}(t+1)\frac{\partial\sigma^{2}\Pi^{-1}(t+1)}{\partial\eta_{k}(t+1)}\hat{\theta}(t+1)\right)$ (42)